OpenStudy (anonymous):

Secant RM intersects secant RN at point R. Find the length of RQ . If necessary, round to the hundredths place.

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):

what does the theorem say ...\[RP \times RN=RQ \times RM\]

OpenStudy (paxpolaris):

\[3\times8 = x \times \left( x+9 \right)\\ \implies x^2+9x-24=0\\ \implies x= {-9 + \sqrt{177} \over 2}\]